OpenStudy (unimatix):

Determine in the following case whether there exists a real number x satisfying the indicated relation, and if there is, determine this number.

OpenStudy (unimatix):

The answer key gives says there is no real number x, but I am having trouble understanding this. \[\sqrt{x-2} = 3 - 2 \sqrt{x}\]

OpenStudy (unimatix):

So what I've tried: \[x - 2 = 9 - 12\sqrt{x} + 4x\]

OpenStudy (unimatix):

then: \[-2 = 9 - 12\sqrt{x} + 3x\] \[-11 = - 12\sqrt{x} + 3x\]

OpenStudy (unimatix):

And then I guess I could divide by - 12? Which gives: \[\frac{ 11 }{ 12 } = \sqrt{x} - \frac{ 1 }{ 4 } x\] But that doesn't really seem to help.

jimthompson5910 (jim_thompson5910):

subtract 3x from both sides of this equation \[\Large -11 = - 12\sqrt{x} + 3x\] then square both sides

OpenStudy (unimatix):

So \[-11-3x = -12\sqrt{x}\] \[121 + 66x + 9x^2= 144x\]

jimthompson5910 (jim_thompson5910):

now get everything to one side and then use the quadratic formula

OpenStudy (unimatix):

So you should get a rational number because you can get a number with the quad. formula?

OpenStudy (unimatix):

actually nope. because number is neg in square. Thank you!

jimthompson5910 (jim_thompson5910):

you should get real number solutions for that last equation you got

jimthompson5910 (jim_thompson5910):

`The answer key gives says there is no real number x` that's not correct. There is one real number solution

OpenStudy (unimatix):

Yeah. I'm getting: \[\frac{ 88\pm \sqrt{3388} }{ 18 }\]

jimthompson5910 (jim_thompson5910):

neither of those are what I'm getting

OpenStudy (unimatix):

Oh. I did some bad arithmetic. Let's see. \[9x^2 -78x + 121 = 0\]

jimthompson5910 (jim_thompson5910):

so far so good

OpenStudy (unimatix):

So \[\frac{ -78 \pm \sqrt{1728} }{ 18 }\]

jimthompson5910 (jim_thompson5910):

nope

OpenStudy (unimatix):

That should be a positive 78

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now you must check each potential solution back into the original equation

jimthompson5910 (jim_thompson5910):

what I would do is find the decimal form of each solution, and then plug them in to test

jimthompson5910 (jim_thompson5910):

example: x = 1/2 is a bit clunky but x = 0.5 is easier to work with

OpenStudy (unimatix):

Okay, so the answer is \[\frac{ 78 - \sqrt{1728} }{ 18 }\]

OpenStudy (unimatix):

Thank you!

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

you're welcome